A 50.0-kg child stands at the rim of a merry-go-round of radius 2.70 m, rotating with an angular speed of 4.00 rad/s. (a) What is the child's centripetal acceleration? (b) What is the minimum force between her feet and the floor of the carousel that is required to keep her in the circular path? N (c) What minimum coefficient of static friction is required? Is the answer you found reasonable? In other words, is she likely to stay on the merry-go-round? Yes No
Answer #1

a] Centripetal acceleration a = w^2 r = 4^2*2.7 = 43.2 m/s^2

b] Minimum force = m*a = 50*43.2 = 2160 N

c] Coefficient of friction u = friction/normal = 2160/(50*9.8) = 4.408

The answer is not reasonable as coefficient of friction required is too high, so she is not likely to stay on the merry-go-round.